Consider the gas-phase reaction 2 Q (g) ⇌ Z (g) at equilibrium in a
container with a moveable piston at 25 °C. The total pressure in the container is 1.1 atm
and PQ = PZ.

1) Calculate Kp for the reaction at 25 °C
2) The piston is moved to change the volume of the container, resulting in a new total
pressure of 1.4 atm. Assuming no change in temperature, calculate the new partial
pressures of Q and Z when equilibrium is restored
1
Expert's answer
2020-03-31T13:42:39-0400

1) "Kp=\frac{Z}{Q^2}"

PQ=PZ

Kp= 1.1/1.12

Kp= 0.5

2) "Kp=\frac{PZ}{PQ^2}"

Kp=1.4/1.42

Kp=0.5

but PZ=PQ

P=0.55